The length of a rectange is 7 times its width. The perimeter of the rectange is 40 feet. What are the dimensions of the rectangle?
Thank you

Answer by LinnW(1048)

W = width
L = length
2W + 2L = 40
L = 7W
2W + 2(7W) = 40
2W + 14W = 40
16W = 40
divide each side by 16
W = 40/16
W = 5/2
L = 7W = 7(5/2) = 35/2
The perimeter is 35/2 + 5/2 + 35/2 + 5/2 = 80/2 = 40
so our values for L and W check out